OpenStudy (anonymous):

Factor completely and then place the factors in the proper location on the grid. x2 - 10x + 24

OpenStudy (unklerhaukus):

can you think of two numbers that add to give -10 and multiply to give 24

OpenStudy (unklerhaukus):

(hint: they are both negative)

OpenStudy (anonymous):

6 and 4

OpenStudy (unklerhaukus):

-4 & -6 add to -10 and multiply to 24 so \[x^2 - 10x + 24=(x-4)(x-6)=x^2-6x-4x+24\]

OpenStudy (unklerhaukus):

so x^2 - 10x + 24 = 0 when x is ether 4 or 6

OpenStudy (unklerhaukus):